In class, I "drew a circuit" with a pencil. When the LED was moderately bright about 9mA was flowing and so (since I was using a 9V battery), the resistance was about 1 kΩ.

At this point, the distance
1
Expert's answer
2019-09-16T09:45:08-0400

From ohm's law,

We know that

"R=\frac{V}{I}"

Or,

"V=RI=10^3\times9\times10^{-3}=9\ V\\So\ potential\ drop across\ LED\ will \ be\ 9\ V"

Power deliverd by LED will be "VI=9\times9\times10^{-3}=8.1\times10^{-2}\ W"

So at this time work done by LED will be "8.1\times10^{-2} \ J\ per\ second"